how can i solve this system of equation by subtraction
4x-y=41
2x+y=25

Answer by stanbon(75887)

solve this system of equation by subtraction
4x-y=41
2x+y=25
---
Add the equations to get rid of the y-terms:
6x = 66
x = 11
---
Now, solve for "y":
2x+y = 25
2*11+y = 25
22 +y = 25
y = 3
============
Cheers,
Stan H.
===========